# Explain Yourself, Briefly! Self-Explaining Neural Networks with Concise Sufficient Reasons

### v2

@@ Appendix A Computational Complexity Background
Appendix B Proof of Theorem 1
@@
Given a neural network classifier $f$ with ReLU activations, and $\textbf{x}\in\mathbb{R}^{n}$, obtaining a cardinally minimal sufficient reason for $\langle f,\textbf{x}\rangle$ is (i) NP-Complete for baseline sufficient reasons (ii) $\Sigma^{P}_{2}$-Complete for robust sufficient reasons and (iii) $\text{NP}^{\text{PP}}$-Hard for probablistic sufficient reasons.
@@
We observe that since the previous proof establishes hardness over the uniform distribution, this hardness persists under any distributional assumption that includes this distribution, such as independent distributions or more expressive ones, like the Markovian distributions examined by (Marzouk & de La Higuera, ).
Appendix C Proof of Theorem ii
@@
Given a neural network classifier $f$ with ReLU activations, and $\textbf{x}\in\mathbb{R}^{n}$, $\forall\epsilon>0$ approximating cardinally minimal probablistic sufficient reasons with $n^{1-\epsilon}$ factor is NP-Hard
@@
Given a CNF classifier, $\psi$, and some instance x, for all $\epsilon>0$ obtaining a probablistic sufficient reason concerning $\langle\psi,\textbf{x}\rangle$ under the uniform distribution defined over $\{0,1\}^{n}$, is NP-Hard to approximate within factor $n^{1-\epsilon}$.
@@
Appendix D Technical Specifications
@@
Image Classification Experiments. All image classification configurations underwent a grid search across various learning rates $\alpha$, with values $\{10^{-2},10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\}$ for both standard and SST training. For SST, additional grid searches were conducted for the cardinality coefficient $\xi$ options: $\{10^{-5},10^{-6},10^{-7},10^{-8},10^{-9},10^{-10},10^{-11}\}$. All models were trained using the Adam optimizer, a batch size of $64$, and held-out validation and test sets. For robust masking, a PGD $\ell_{\infty}$ attack with $\epsilon:=0.12$ was used, consisting of $10$ steps each of size $\alpha^{\prime}:=10^{-2}$. The value of $\epsilon:=0.12$ was selected to strike a balance between allowing a sufficiently large perturbation for learning, while avoiding excessive distortion of the image. On one hand, this creates a challenging task for learning (as opposed to smaller perturbations). On the other hand, it prevents excessive distortion (as might occur with larger perturbations), which could result in the complementary set $\overline{S}$ containing images that are entirely outside the desired distribution. Lastly, we apply a black background to mask the MNIST images, both for SST and post-hoc approaches, leveraging their naturally zeroed-out borders, which provides a clearer interpretation of the portions of the digit that fall within the sufficient reason. Moreover, for CIFAR-10 and IMAGENET images, we select images from these benchmarks where a specific object is highlighted against a white background. The sufficient reasons for these images are depicted over the same white background, allowing the explanation to remain focused on the object itself. This approach is applied consistently for both SST and the post-hoc methods.
Language Classification Experiments. All language classification experiments utilized a pre-trained Bert-base (Devlin et al. (2018)) model (applicable to both standard training and SST). The grid search focused on learning rates $\alpha:=\{2e^{-5},3e^{-5},5e^{-5}\}$, which are the typical values used for optimizing pre-trained Bert models (Devlin et al. (2018)). For SST, an additional grid search was conducted for the cardinality coefficient $\xi$ options: $\{10^{-4},10^{-5},10^{-6},10^{-7},10^{-8}\}$. Optimization was performed using the standard AdamW optimizer, with a batch size of $32$ and held-out validation and test sets.
D.1 MNIST
@@ D.5 IMDB
Appendix E Training Time
@@
Image classification tasks. For MNIST, Standard training required $311.03$ seconds (over 43 epochs), whereas baseline-masking SST was completed in just $26.55$ seconds (over 2 epochs). This significant speed increase in the training, despite involving a dual-propagation process, was primarily due to the optimal learning rate for SST with baseline masking being set at $10^{-3}$, compared to $10^{-4}$ for standard training. For probabilistic masking, the training duration was $180.83$ seconds (over 11 epochs), where the enhanced efficiency stemmed, again, from achieving optimal results at a higher learning rate. However, robust masking, which is more computationally demanding, took substantially longer, clocking in at $1796.99$ seconds (over 49 epochs). For IMAGENET, standard training using the robust masking configuration ran for $287056.81$ seconds (over $22$ epochs) while the parallel standard-training configuration for IMAGENET ran for $74141.85$ seconds, over $43$ epochs (around 4 times faster). For CIFAR-10, SST using the robust masking that was mentioned in the paper took $3159.15$ seconds (over 53 epochs) compared to standard training which took 477.42 seconds (over 49 epochs).
@@
Language Classification tasks. For IMDB sentiment analysis, standard training ran for $4974.33$ seconds (over $1$ epoch), while SST with baseline masking ran for $9713.42$ seconds (over $2$ epochs) and SST with probabilistic masking clocked at $21516.37$ seconds (over $3$ epochs). For SNLI classification, standard training ran for $5353.71$ seconds (over $2$ epochs), while SST with a probabilistic masking ran for $6084.66$ seconds (over $1$ epoch), SST with a baseline masking ran for $10440$ seconds (over $2$ epochs).
Appendix F Supplementary Results

### v3

@@ Appendix B Computational Complexity Background
Appendix C Proof of Theorem 1
@@
Given a neural network classifier $f$ with ReLU activations, and $\textbf{x}\in\mathbb{R}^{n}$, obtaining a cardinally minimal sufficient reason for $\langle f,\textbf{x}\rangle$ is (i) NP-Complete for baseline sufficient reasons (ii) $\Sigma^{P}_{2}$-Complete for robust sufficient reasons and (iii) $\text{NP}^{\text{PP}}$-Hard for probabilistic sufficient reasons.
@@
We observe that since the previous proof establishes hardness over the uniform distribution, this hardness persists under any distributional assumption that includes this distribution, such as independent distributions or more expressive ones, like the Markovian distributions examined by (Marzouk & de La Higuera, ), or the Hidden Markov modeled distributions examined by (Marzouk et al., 2025).
Appendix D Proof of Theorem ii
@@
Given a neural network classifier $f$ with ReLU activations, and $\textbf{x}\in\mathbb{R}^{n}$, $\forall\epsilon>0$ approximating cardinally minimal probabilistic sufficient reasons with $n^{1-\epsilon}$ factor is NP-Hard
@@
Given a CNF classifier, $\psi$, and some instance x, for all $\epsilon>0$ obtaining a probabilistic sufficient reason concerning $\langle\psi,\textbf{x}\rangle$ under the uniform distribution defined over $\{0,1\}^{n}$, is NP-Hard to approximate within factor $n^{1-\epsilon}$.
@@
Appendix E Technical Specifications
@@
E.1 Image Classification Experiments
All image classification configurations underwent a grid search across various learning rates $\alpha$, with values $\{10^{-2},10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\}$ for both standard and SST training. For SST, additional grid searches were conducted for the cardinality coefficient $\xi$ options: $\{10^{-5},10^{-6},10^{-7},10^{-8},10^{-9},10^{-10},10^{-11}\}$. All models were trained using the Adam optimizer, a batch size of $64$, and held-out validation and test sets. For robust masking, a PGD $\ell_{\infty}$ attack with $\epsilon:=0.12$ was used, consisting of $10$ steps each of size $\alpha^{\prime}:=10^{-2}$. The value of $\epsilon:=0.12$ was selected to strike a balance between allowing a sufficiently large perturbation for learning, while avoiding excessive distortion of the image. On one hand, this creates a challenging task for learning (as opposed to smaller perturbations). On the other hand, it prevents excessive distortion (as might occur with larger perturbations), which could result in the complementary set $\overline{S}$ containing images that are entirely outside the desired distribution. Lastly, we apply a black background to mask the MNIST images, both for SST and post-hoc approaches, leveraging their naturally zeroed-out borders, which provides a clearer interpretation of the portions of the digit that fall within the sufficient reason. Moreover, for CIFAR-10 and IMAGENET images, we select images from these benchmarks where a specific object is highlighted against a white background. The sufficient reasons for these images are depicted over the same white background, allowing the explanation to remain focused on the object itself. This approach is applied consistently for both SST and the post-hoc methods.
E.2 Language Classification Experiments
All language classification experiments utilized a pre-trained Bert-base (Devlin et al. (2018)) model (applicable to both standard training and SST). The grid search focused on learning rates $\alpha:=\{2e^{-5},3e^{-5},5e^{-5}\}$, which are the typical values used for optimizing pre-trained Bert models (Devlin et al. (2018)). For SST, an additional grid search was conducted for the cardinality coefficient $\xi$ options: $\{10^{-4},10^{-5},10^{-6},10^{-7},10^{-8}\}$. Optimization was performed using the standard AdamW optimizer, with a batch size of $32$ and held-out validation and test sets.
E.3 MNIST
@@ E.7 IMDB
Appendix F Training Time
@@ F.1 Image classification tasks
For MNIST, Standard training required $311.03$ seconds (over 43 epochs), whereas baseline-masking SST was completed in just $26.55$ seconds (over 2 epochs). This significant speed increase in the training, despite involving a dual-propagation process, was primarily due to the optimal learning rate for SST with baseline masking being set at $10^{-3}$, compared to $10^{-4}$ for standard training. For probabilistic masking, the training duration was $180.83$ seconds (over 11 epochs), where the enhanced efficiency stemmed, again, from achieving optimal results at a higher learning rate. However, robust masking, which is more computationally demanding, took substantially longer, clocking in at $1796.99$ seconds (over 49 epochs). For IMAGENET, standard training using the robust masking configuration ran for $287056.81$ seconds (over $22$ epochs) while the parallel standard-training configuration for IMAGENET ran for $74141.85$ seconds, over $43$ epochs (around 4 times faster). For CIFAR-10, SST using the robust masking that was mentioned in the paper took $3159.15$ seconds (over 53 epochs) compared to standard training which took 477.42 seconds (over 49 epochs).
F.2 Language Classification tasks
@@
For IMDB sentiment analysis, standard training ran for $4974.33$ seconds (over $1$ epoch), while SST with baseline masking ran for $9713.42$ seconds (over $2$ epochs) and SST with probabilistic masking clocked at $21516.37$ seconds (over $3$ epochs). For SNLI classification, standard training ran for $5353.71$ seconds (over $2$ epochs), while SST with a probabilistic masking ran for $6084.66$ seconds (over $1$ epoch), SST with a baseline masking ran for $10440$ seconds (over $2$ epochs).
Appendix G Generalization between Sufficiency Settings
In this section, we examine SST’s ability to generalize across different types of sufficiency configurations. Our work considers various sufficiency forms, including baseline, probabilistic, and robust masking configurations. Users can select the masking procedure that aligns with their preferred sufficiency criterion. Stricter sufficiency constraints typically lead the model to select larger subsets, while looser constraints result in smaller ones. However, training SST with a specific masking configuration for one form of sufficiency does not necessarily ensure strong generalization to other forms. A potential approach to improve generalization is to train SST with multiple masking configurations, such as varying the masking setting across batches — an avenue worth exploring in future work. Nevertheless, even when trained with a single masking configuration, there are cases where one sufficiency form may implicitly imply another, particularly when it encompasses other forms. For instance, uniform sampling within a bounded perturbation may be naturally subsumed by a robust masking strategy applied to the same perturbation.
We conducted an additional experiment in which we trained each model from the experiments in Tables 2 and 3 (covering MNIST, SNLI, and IMDB) to evaluate the impact of different masking configurations on various forms of sufficiency. The results are presented in Table 4.
Table 4: The generalization of various masking configurations to different sufficiency conditions.
Masking Faithfulness
  Robust  Probabilistic Baseline
robust 99.28 99.32 11.82
MNIST baseline 98.91 98.38 96.52
probabilistic 98.85 99.11 8.16
SNLI baseline 44.81 95.88
probabilistic 95.35 93.12
IMDB baseline 75.7 98.05
probabilistic 95.67 77.7
Table 4 presents results highlighting key generalization patterns across different tasks. For language tasks, we observe the following: In SNLI, probabilistic masking demonstrates high baseline faithfulness, whereas baseline masking exhibits considerably lower probabilistic faithfulness, suggesting superior generalization under probabilistic constraints. In IMDB, both probabilistic and baseline masking achieve moderate faithfulness in the opposite masking configuration. In the MNIST experiment, a more intricate dynamic emerges. Robust masking and probabilistic sampling remain within a bounded $\epsilon_{p}$ domain, while the baseline configuration is significantly out-of-distribution (OOD), making the baseline task more challenging. This is reflected in the larger subset size (23.69%) produced by SST for the baseline compared to the robust and probabilistic settings. As a result, the larger sufficient reason derived from the baseline generalizes well across other configurations, maintaining high faithfulness. Conversely, subsets from probabilistic and robust masking generalize effectively to each other but perform poorly on the baseline, likely due to the baseline’s pronounced OOD nature. These findings emphasize the complex interaction between domain selection, sampling distribution, and baseline properties in influencing generalization. We believe further exploration into the generalization capabilities of SST across masks presents an exciting research direction.
G.1 Discussion on variations of the faithfulness metric
Similar to the challenges that arise when generalizing across different sufficiency configurations, a similar concern extends to the evaluation of explanations — specifically, the computation of faithfulness. As with many explanation frameworks, defining a sufficient reason can be ambiguous, particularly due to the varying interpretations of “missingness” in the complement. Our work addresses this inherent limitation, which is prevalent in many post-hoc methods, by evaluating and training explanations across diverse sufficiency formulations. However, we recognize that no single definition can fully encapsulate a model’s internal behavior, and this limitation also applies to faithfulness metrics assessing the sufficiency of subsets. This issue is widespread in many XAI methods, including SHAP (Sundararajan & Najmi, 2020), and metrics such as infidelity (Yeh et al., 2019), where different treatments of “missing” features lead to variations in explanation definitions and evaluation metrics. By exploring diverse sufficiency criteria, our work contributes to a more comprehensive understanding of this class of explanations.
Appendix H Extensions to Simplified Input Spaces
It is evident that SST can be trained using simplified input spaces instead of individual features such as pixels or tokens. These simplified spaces — such as super-pixels, sentences, or other meaningful units — can, in some cases, enhance the human interpretability of the resulting explanations, as explored in previous works (e.g., (Ribeiro et al., 2016; 2018)). To evaluate this approach, we conducted an experiment on MNIST and CIFAR-10, using a straightforward simplified input setting of $2\times 2$ pixel patches. Accordingly, we adjusted the output vector to match the number of super-pixels and applied masking at the level of these simplified inputs.
For consistency, we used the same hyperparameter choices as in the original (non-simplified input space) grid search. However, a dedicated grid search for this specific configuration could yield improved results. Our findings indicate that, for MNIST, the average explanation size was 4.26%, with 90.53% faithfulness and a minimal accuracy drop of -0.34%. In contrast, for CIFAR-10, we observed a -4.54% accuracy drop, 91.74% faithfulness, and an explanation size of 22.93%. These results suggest that, under this particular setting, the original (non-simplified) input space produced better overall performance metrics. However, from a human interpretability standpoint, simplified inputs can provide more intuitive and comprehensible explanations. We include examples of these simplified explanations alongside their regular counterparts in Figure 6.
Figure 6: Examples of comparisons between explanations generated by SST using superpixels with a $2\times 2$ patch versus those based on individual pixels.
Appendix I Additional Ablation Experiments
In this section, we provide additional ablation results for our experiments. Specifically, we examine the effects of varying both the $\epsilon_{p}$ perturbation and the $\alpha$ step size in the robust masking experiments, as well as adjusting the $\tau$ selection threshold parameter. The results are summarized in Tables 6, 6 and 7, with a detailed discussion of each experiment provided separately.
I.1 Varying the $\epsilon_{p}$-bound in robust masking
We perform an ablation study by varying the $\epsilon_{p}$ attack in robust masking on CIFAR-10, using the same configurations as in our main experiment. In general, increasing $\epsilon_{p}$ makes it more challenging to satisfy the sufficiency conditions of the explanation. Consequently, users can select different values based on the desired "degree" of sufficiency. A higher $\epsilon_{p}$ can lead to larger explanation sizes and/or reduced faithfulness, or it may result in decreased model accuracy. The specific metric affected depends on the relative weighting of different loss coefficients during training, which influence the optimization objectives.
As shown in Table 6, increasing $\epsilon_{p}$ within our configurations generally leads to lower faithfulness and larger explanation sizes. However, we note that excessively large values of $\epsilon_{p}$ can reduce the effectiveness of gradient perturbations, a well-known issue in adversarial training (Addepalli et al., 2022). Thus, it is not surprising that at very high perturbations (e.g., above $0.25$), model accuracy deteriorates significantly. Addressing the challenges posed by such large perturbations, which is also relevant to adversarial training, remains an open problem for future research.
I.2 Varying the step size $\alpha$ in robust masking.
We conduct an ablation study by adjusting the step size $\alpha$ in adversarial training on CIFAR-10, while keeping the adversarial perturbation $\epsilon_{p}$ fixed at $0.12$. The results are presented in Table 6. Similar to adversarial training, smaller values of $\alpha$ yield adversarial examples that are closer to the original input, whereas larger values generate more distant examples. Importantly, the faithfulness measure we present quantifies the proportion of test instances that remain sufficient under a robust mask for each configuration, all computed using the same attack strategy (i.e., with the same $\alpha$ value). These findings demonstrate that our model can adapt to different $\alpha$ values, producing relatively compact and faithful explanations. However, a substantial increase in step size leads to a decline in accuracy.
Table 5: An ablation study examining the effect of varying $\epsilon_{p}$ perturbations in robust masking on the CIFAR-10 benchmark.
$\epsilon_{p}$ (%) Acc. (%) Faith. (%) Size (%)
0.01 -0.13 99.31 10.2
0.05 +0.7 95.09 0.91
0.1 -1.58 91.46 9.78
0.12 -1.96 90.43 12.99
0.15 -1.45 88.03 11.91
0.2 -3.32 85.34 25.49
0.25 -4.58 79.9 17.48
0.3 -9.92 78.88 35.9
0.35 -5.97 70.94 19.2
0.4 -10.25 64.95 34.53
0.45 -14.68 64.95 34.53
0.5 -14.09 68.05 31.99
Table 6: An ablation study examining the effect of varying $\alpha$: the step size in robust masking (with a fixed $\epsilon_{p}:=0.12$ perturbation) on the CIFAR-10 benchmark.
$\alpha$ (%) Acc. (%) Faith. (%) Size (%)
1/255 -0.66 89.47 12.01
2/255 -2.22 89.9 19.7
3/255 -1.13 89.74 10.93
4/255 -2.12 90.83 17.28
5/255 -1.96 90.43 12.99
6/255 -0.08 89.25 7.2
7/255 -1.32 90.21 9.78
8/255 -1.91 90.44 15.09
9/255 -1.68 90.27 12.89
10/255 -2.71 89.15 17.27
Table 7: An ablation study examining the effect of varying the selection threshold $\tau$ in probabilistic masking on the CIFAR-10 benchmark.
$\tau$ (0-1 range) Acc. Gain (%) Faith. (%) Size (%)
0.1 +0.39 73.88 3.96
0.2 -0.61 81.22 5.81
0.3 -0.23 81.26 2.55
0.4 -1.57 84.68 9.80
0.5 -0.09 87.54 9.08
0.6 -1.62 88.32 12.26
0.7 -1.79 96.44 11.80
0.8 -2.08 86.97 12.63
0.9 -1.24 86.43 6.65
I.3 Varying the selection threshold $\tau$
Lastly, we performed an ablation study on different values of $\tau$ trained with probabilistic masking on CIFAR-10. The results are presented in Table 7. Overall, as observed in the results, our method adapts to varying values of $\tau$ by optimizing the model to learn different weights within the explanation component output vector. However, excessively high or low $\tau$ values tend to be more sensitive and may negatively impact certain key metrics. For instance, a low setting of $\tau:=0.1$ resulted in reduced faithfulness ($73.88\%$), while a high setting of $\tau:=0.8$ led to lower accuracy ($-2.08\%$). In contrast, setting $\tau:=0.5$ yielded comparable accuracy ($-0.09\%$).
Appendix J Supplementary Results